OpenStudy (anonymous):

HELP i will meadel and Fan!!!!!!!! the cost of a student ticket to the school play is 7$. write an equation that correctly relates the cost, c, for a particular number of tickets, purchased. identify the independent and dependent variable.

Mehek (mehek14):

y = 7x

Mehek (mehek14):

x is independent y is dependent on x

OpenStudy (shamim):

Let the number of ticket=x

OpenStudy (anonymous):

the number of the student tickets are s so what would it be

OpenStudy (shamim):

Total cost for x ticket=7x

OpenStudy (anonymous):

oka what is the independent varible and the dependent varible

OpenStudy (shamim):

So c=7x Right?

OpenStudy (anonymous):

i dont know that is why i asked u guys

OpenStudy (shamim):

x is independent variable

OpenStudy (shamim):

c is dependent variable

OpenStudy (igreen):

@Mehek14 Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. Please read the \(\bf\color{blue}{\underline{{\href{http:///openstudy.com/users/mlpsisters7}{Code-of-Conduct}}}}\) to learn more.

OpenStudy (anonymous):

i dont understan i reen can you explain it to me more please

OpenStudy (anonymous):

so x representss what

OpenStudy (shamim):

If x=2 Then c=2*7=14 Right?

OpenStudy (shamim):

And if x=5 Then y=5*7=35 Right?

OpenStudy (shamim):

So if the value of x changes then the value of c will b changes

OpenStudy (shamim):

So the value of c will depend on the value of x

OpenStudy (shamim):

I think u got it?

OpenStudy (shamim):

x represents the number of tickets